### v1

# Multi-state Protein Design with DynamicMPNN

Abstract
Structural biology has long been dominated by the one sequence, one structure, one function paradigm, yet many critical biological processes—from enzyme catalysis to membrane transport—depend on proteins that adopt multiple conformational states. Existing multi-state design approaches rely on post-hoc aggregation of single-state predictions, achieving poor experimental success rates compared to single-state design. We introduce DynamicMPNN, an inverse folding model explicitly trained to generate sequences compatible with multiple conformations through joint learning across conformational ensembles. Trained on 46,033 conformational pairs covering 75% of CATH superfamilies and evaluated using AlphaFold initial guess, DynamicMPNN outperforms ProteinMPNN by up to 13% on structure-normalized RMSD across our challenging multi-state protein benchmark.
1 Introduction
The conformational diversity of proteins underlies biological functions such as enzyme catalysis, protein recognition, allostery, protein evolution modulation or human disease (Monzon et al., 2016). Of particular importance is the design of bioswitches, with huge implications in biotechnology, e.g. in creating artificial motors, signalling pathways, biosensors or drug delivery systems (Stein & Alexandrov, 2015; Praetorius et al., 2023). While most known switches undergo rearrangements in the context of a single fold (Ambroggio & Kuhlman, 2006a), the class of metamorphic proteins undergo changes in their secondary structure and fold (Fig. 1a) and have been predicted to represent up to 4% of the PDB chains (Porter & Looger, 2018). These proteins are known to take two main functional states (Dishman & Volkman, 2018) and a finite number of conformations (see Discussion), thus this study is not aimed at the design of protein sequences adopting continuous conformational landscapes (e.g. intrinsically disordered proteins (Tompa & Fuxreiter, 2008)).
Multi-state protein design was first achieved through rational design and physics-based RosettaDesign (Liu & Kuhlman, 2006) of: metamorphic metal-binding peptides (Ambroggio & Kuhlman, 2006b; Cerasoli et al., 2005); closely related sequences that adopt diverging folds (Wei et al., 2020); and hinge proteins with binder-regulated equilibrium between states (Zhang et al., 2022; Quijano-Rubio et al., 2021; Praetorius et al., 2023). Remarkably, Praetorius et al. (2023) additionally employed ProteinMPNN Multi-state Design (ProteinMPNN-MSD) (Dauparas et al., 2022)—an inference strategy for extending ProteinMPNN to multiple states by averaging the logits of two independent single-state ProteinMPNN embeddings during the decoding step—to design proof-of-concept multi-helix hinge proteins.
Praetorius et al. (2023) however report an overall very low experimental success rate: only 46 2-state hinge sequences, corresponding to roughly 0.002% of all designs, were successfully expressed in solution out of which only nine showed binding with the corresponding target peptide. This and the absence of any reported de novo designs for full natural multi-state backbones is proof that multi-state protein design lags behind single-state design (Appendix E) due to limited conformational datasets, weak benchmarks, and the poor record of folding models in predicting alternative states (Chakravarty et al., 2024) - which made for poor self-consistency filters.
Our contributions. This paper introduces DynamicMPNN (Fig. 1b), a novel geometric deep learning-based pipeline for multi-state protein sequence design.
DynamicMPNN is the first explicit multi-state inverse folding model for protein design. To train DynamicMPNN, we create a new ML-ready dataset of proteins with multiple conformations using the PDB and CoDNaS (Monzon et al., 2016) databases, and evaluate the method on 94 biologically relevant metamorphic, hinge, and transporter proteins.
DynamicMPNN improves performance over ProteinMPNN (Dauparas et al., 2022) on AFIG by up to 13% on RMSD and 3% on pLDDT self-consistency values.
2 The DynamicMPNN pipeline
2.1 Protein multi-conformational dataset
While over 750,000 individual protein chains (sequence-structure pairs) are available in the PDB, multi-conformational data is far more scarce with only 11,833 NMR-derived protein ensembles covering just 21% of CATH superfamilies. To overcome this limitation, we exploit the redundancy of chains with very high sequence similarity (95%) in the PDB to build a multi-conformational dataset of 46,033 conformer pairs that expands coverage to 75% of CATH superfamilies (Fig. 2a).
For dataset splitting, we first curate a benchmark set from five previous studies of proteins with large conformational changes: 92 metamorphic proteins (Porter & Looger, 2018), 91 apo-holo proteins (Saldaño et al., 2022), the OC23/OC85 open-closed datasets, and 20 transporter proteins (Kalakoti & Wallner, 2025). The 94 highest RMSD pairs were assigned to the test set and the next 100 to the validation set. Training clusters were filtered to exclude any with TM-score (Zhang & Skolnick, 2004) to test/validation structures, preventing structural similarity leakage and yielding a final training set of 44,243 conformer pairs. See Appendix B for further details on dataset composition.
2.2 DynamicMPNN for multi-state inverse folding
Single-state inverse folding methods seek to model the conditional distribution where represents a protein backbone with residues, and is the amino acid sequence. Extensions of these methods to multi-state design have thus far been limited to post-hoc aggregation of independent single-state predictions. Instead, DynamicMPNN learns the joint conditional distribution of directly through autoregressive sequence generation (see Appendix C), where represent distinct conformations; thus capturing sequence compatibility across all states simultaneously within a single model.
Overall architecture. DynamicMPNN independently encodes each of the functional states of the protein, together with their interaction partners, into a shared latent feature space (Fig. 1b). Embeddings of the target chains are then pooled across conformations to obtain a single embedding from which a sequence is auto-regressively generated.
Our architecture is based on gRNAde (Joshi et al., 2025), a multi-state GNN for RNA inverse folding. For both the encoder and the decoder, we employ SE(3)-equivariant Geometric Vector Perception (Jing et al., 2021) layers which maintain computational efficiency through edge sparsity. Both encoder and decoder were assigned 8 GVP layers, following Hsu et al. (2022). See Appendix D for further details.
Alignment and pooling. To deal with nonidentical sequences and missing residues during training, structure pairs are featurized and encoded independently, and subsequently aggregated based on pairwise sequence alignments (i.e., only nongap residues are taken into account when pooling).
Multi-chain encoding and masking. Given that conformational shifts often depend on interactions, DynamicMPNN encodes the chemical environment for each conformational state (currently limited to proteins). During training, we expose the sequence information of binding partners in the encoding, only masking chains with 70% sequence similarity to the ground truth sequence of the chain of interest.
Conformation order-invariant pooling. We employ Deep Set pooling (Zaheer et al., 2017) over the conformations as it is invariant to conformation order and does not add extra parameters to our model. We note that while some more expressive pooling strategies have been shown to provide marginal performance improvements, they usually come at a great cost in efficiency (Joshi et al., 2025).
| RMSD | pLDDT | |||||||
| Model | Best Paired | Best Paired | Best Paired | Best Single | All Avg. | Best Paired | Best Single | All Avg. |
| (Å) | (Struct Norm) | (Decoy Norm) | (Å) | (Å) | ||||
| DynamicMPNN | 13.43 | 3.86 | 0.58 | 12.76 | 17.39 | 58.65 | 59.31 | 50.22 |
| ProteinMPNN-MSD | 14.76 | 4.47 | 0.65 | 13.99 | 17.24 | 55.89 | 57.74 | 50.18 |
| Natural sequence | 16.85 | 5.51 | 0.76 | 16.85 | 16.85 | 48.51 | 48.51 | 48.51 |
2.3 Multi-state design evaluation
Following previous work (Wang et al., 2023), we evaluate the refoldability of generated sequences rather than sequence recovery (Appendix A). Existing refoldability methods compare target structures to single conformations predicted by folding models (e.g., AlphaFold2 (Jumper et al., 2021)). We argue that this approach is unsuitable for multi-state design since folding models typically predict one dominant state or interpolate between conformations rather than sampling the full conformational ensemble (Lane, 2023; Chakravarty et al., 2024; Saldaño et al., 2022).
We propose using the AlphaFold initial guess (AFIG) framework (Roney & Ovchinnikov, 2022; Bennett et al., 2023), originally suggested to rank candidate protein structures, to evaluate multi-state designs (Praetorius et al., 2023). AFIG initializes the AF2 reference frames on the target backbone coordinates during inference to bias the model towards the target conformation. The similarity (C-RMSD) between predicted and target structures, along with AF2 confidence scores, serves as a proxy for the likelihood that the designed sequence will fold into the target structure. This framework naturally enables evaluation of refoldability across multiple conformations.
Formally, for a protein with conformational states and designed sequence , we define the AFIG RMSD for each target conformation as:
where is the structure predicted by AlphaFold2 for sequence when backbone frames are initialized on coordinates of structure . We define two normalization criteria to contextualize observed deviations:
Structure normalization (Struct Norm): We normalize AFIG RMSD by the maximum RMSD between target conformations to capture task difficulty; i.e. designing proteins with large conformational shifts is inherently more difficult than small rearrangements:
Decoy normalization (Decoy Norm): We initialize AF2 with structurally dissimilar decoy structures () using the same sequences designed and measure the resulting deviations. This control assesses whether sequences fold specifically into their targets or may fold equally well into arbitrary structures:
where is a decoy that is structurally dissimilar to .
Additionally, we measure pLDDT confidence scores to evaluate AFIG fold uncertainty. High RMSD with low pLDDT indicates poor template matching, while low RMSD with high pLDDT suggests a successful design.
3 Results and Discussion
Setup. DynamicMPNN was trained for 50 epochs, and an AFIG evaluation on the validation set was performed every third epoch to select the best-performing model. Then, DynamicMPNN and ProteinMPNN (using Multi-state Design inference strategy) were used to sample 16 sequences for the benchmark set of 94 paired conformations, which were each run through the AFIG pipeline separately against both target states.
Aggregated evaluation metrics were computed as follows. (Best Single) The minimum RMSD and maximum pLDDT values across all sets of 16 sequences are selected from among both states. (Best Paired) The RMSD and pLDDT values are averaged over the 2 states for each sequence, then the best RMSD and pLDDT across all 16 averaged values are selected. Struct Norm and Decoy Norm RMSD values are then calculated for each protein. (All Avg.) The RMSD and pLDDT values are averaged across all 16 sequences and both states. The natural sequences were processed identically and resulting metrics were averaged over the benchmark set and reported in Table 1.
DynamicMPNN outperforms ProteinMPNN. We see this across all metrics, with the largest gains in the ”Best Paired” metrics of up to 13% - arguably the most important set of metrics for experimental design, as only the best sequences are selected for in vitro validation. Visual structural inspection of both states (Fig. 5) confirms successful design of the metamorphic Switch Arc protein by DynamicMPNN, compared to failure to recapitulate the central beta sheet fold by the best scoring ProteinMPNN sequence. DynamicMPNN also outperforms ProteinMPNN in the ”Best Single” metrics, which is surprising since ProteinMPNN was trained on single-state design tasks. It should be noted that ProteinMPNN training dataset contains proteins within the sequence clusters of 84 out of 94 benchmark proteins. This represents a substantial data leakage advantage for ProteinMPNN, while DynamicMPNN was specifically trained to exclude any sequence similarity to the test set. We plan on training a single-state version of DynamicMPNN to more fairly demonstrate the advantage of the multi-state design approach. While DynamicMPNN AFIG RMSD values are still high, it is important to contextualize these results within our deliberately challenging benchmark: our test set comprises proteins with the largest documented conformational changes in the known proteome. This can be seen as an extrapolation beyond the overall modest conformational changes observed in most PDB structures and hence our training data.
Natural sequence vs. multi-state designs. While the best-of-16 sequence from DynamicMPNN and ProteinMPNN (Dauparas et al., 2022) outperform natural sequences, the average metrics are slightly worse than the natural sequences (Fig. LABEL:fig:supplementary_plots). Dauparas et al. (2022) have presented a similar finding with respect to single-state protein design, where the ProteinMPNN designed sequences showed improved AlphaFold refoldability over the natural sequences. A possible explanation is that while both inverse folding models are trained mostly to optimize the sequence for structure stability, evolution has undergone a multi-objective optimization for different functions, which might have prevented it from reaching the energy minima for those structures. Additionally, due to graph sparsity, DynamicMPNN and ProteinMPNN emphasize local interactions, making designed sequences more foldable by AFIG - since AFIG does not use any MSA, it poorly captures long-range interactions present in natural sequences (Zhang et al., 2024).
4 Conclusion
We present DynamicMPNN, the first explicit multi-state inverse folding model, achieving up to 13% improvement over ProteinMPNN on our multi-state benchmark. By jointly learning across conformational ensembles rather than aggregating single-state predictions, DynamicMPNN captures sequence constraints required for multiple functional conformations. This opens possibilities for engineering synthetic bioswitches, allosteric regulators, and molecular machines.
Software and Data
We provide the code for our model in the following repository: github.com/Alex-Abrudan/DynamicMPNN.
Acknowledgements
The authors would like to thank Adam Broerman, Rohith Krishna, Kai Yi, Simon Mathis, Hannah Han, Vladimir Radenkovic, Gustavo Parisi, Sebastian Ahnert, and Pietro Liò for insightful comments and discussions. CKJ was supported by the A*STAR Singapore National Science Scholarship (PhD). AK was supported by the Graduate School of Quantitative Biosciences Munich (QBM).
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Supplementary Results
A.1 Additional Performance Analysis
For our statistical analysis, we first evaluated the normality of the paired differences between model performances (RMSD values) using the Shapiro-Wilk test. The test results (W = 0.913, p 0.001) strongly rejected the null hypothesis of normality for the distribution of differences.
Therefore, we selected the Wilcoxon signed-rank test as the appropriate non-parametric alternative to the paired t-test for assessing statistical significance between model performances. This test makes no assumptions about the underlying distribution shape, making it robust for our analysis of paired protein structure comparisons across the different design approaches.
Fig. LABEL:fig:supplementary_plots (a), (e) shows how DynamicMPNN maintains its superiority for the single-state best sequence design for the RMSD and pLDDT metrics. However, averaging the performance of all designed sequences erases the advantage of DynamicMPNN over ProteinMPNN and natural sequences (Fig. LABEL:fig:supplementary_plots (b), (c), (d), (f)).
While ProteinMPNN outperforms DynamicMPNN on sequence recovery and perplexity (Table 2), Wang et al. (2023) argue that refoldability - emphasised in the present work - is more direct, structure-grounded, and superior to sequence recovery and perplexity when it comes to evaluating inverse folding models.
| Model | Sequence Recovery | Perplexity |
| DynamicMPNN | 0.29 | 10.2 |
| ProteinMPNN MSD | 0.38 | 5.12 |
Appendix B Dataset Details
To construct the dataset, we obtained 46,033 Multiple Sequence Alignment (MSA) clusters at local sequence similarity from the latest version of CoDNaS (v2025) (Monzon et al., 2016), including NMR model structures. Importantly, the CoDNaS dataset prevents potential errors created by mixing different homologues in the same cluster by enforcing the same UniProt ID for all custer members. The 95% similarity threshold accounts for soluble tags or point mutations/alterations performed in different experiments so that all available conformations of a protein are included - as different experimental conditions and sequence variations can reveal distinct thermodynamic states of the same protein (Best et al., 2006). While clusters contain varying numbers of conformations (Fig. 2b) we constructed our dataset using only pairs of chains from one or two PDB entries that have the largest RMSD from each cluster (Fig. 2c). Since the proteins targeted by this project explore two functional states, this choice aims to minimize alignment artifacts and maximize conformational signal during training - chosen pairs represent the most distinct conformational states. We will include conformational information of the whole cluster in future work.
Appendix C Multi-conformational inverse folding
Given a set of target conformations we seek to model the conditional probability distribution over amino acid sequences that can adopt all specified structures. The challenge lies in learning a sequence distribution that simultaneously satisfies multiple structural constraints. We decompose this joint conditional probability using the autoregressive factorization:
where each factor represents the probability of selecting residue given the sequence prefix and the complete structural ensemble.
Appendix D Model Details
D.1 Featurisation scheme
We use a similar featurisation scheme as in (Jamasb et al., 2024). Node scalar features are transformer-like positional encoding in a 16-dimensional array; backbone dihedral angles ; the virtual torsion and virtual bond angle . Node vector features are position vectors of , . Edge scalar features are established via k-NN (k=16) and the edge length expressed in 32 Radial Basis Functions, , as well as the length of the edge itself. Edge vector features are edge directional unit vectors for both directions . To further prevent overfitting on crystallisation artifacts, random Gaussian noise (Å) was added to the coordinates (Dauparas et al., 2022).
D.2 Multi-state GNN
DynamicMPNN processes one or multiple protein backbone graphs via a multi-state GNN encoder (Joshi et al., 2025). Overall, DynamicMPNN’s encoder is equivariant to 3D roto-translation of coordinates as well as ordering of the states in its input. Encoding is followed by pooling node features across states, which is invariant to the ordering of the states, and autoregressive sequence decoding.
When representing conformational ensembles as a multi-graph, each node feature tensor contains three axes: (#nodes, #conformations, feature channels). Multi-state GNN’s encode multi-graphs by performing message passing on the multi-graph adjacency to independently process each conformer, while maintaining permutation equivariance of the updated feature tensors along both the first (#nodes) and second (#conformations) axes.
D.3 Geometric Vector Perceptron layers
Geometric Vector Perceptrons (GVPs) (Jing et al., 2021) are a generalization of MLPs to take tuples of scalar and vector features as input and apply -equivariant non-linear updates. GVP GNN layers process scalar and vector features on separate channels to maintain equivariance. The node scalars , node vectors , and edge scalars and vectors communicate through a message passing operation:
| (Message & aggregate steps) | (5) | ||||
| (Update step) | (6) |
The overall GNN encoder is -equivariant due to the use of reflection-sensitive input features (dihedral angles) combined with -equivariant GVP-GNN layers.
D.4 Conformation order-invariant pooling
After using message passing layers that are conformation order-equivariant, we add a conformation order-invariant head, which performs average pooling across the conformation channel of the scalar and vector feature tensors, similar to Joshi et al. (2025): and to and , where is the sequence length, is the number of backbones, is the number of scalar features, and is the number of vector features. The only pooling strategy used in this work is the pooling of the maximum RMSD pair of chains - therefore - although more pooling strategies for homo-oligomers can be used, such as equal averaging of all chains to be inverse folded in the selected PDB entries.
Appendix E Further Discussion
E.1 Motivation
A commonly derived assumption from Anfinsen’s experiment is that proteins adopt only one native 3D structure, leading to the “one sequence, one structure, one function” canon. This view has been indirectly reinforced by the predominant use of X-Ray crystallography in experimental protein structure determination - it requires the protein to form a single, diffractible crystal (Dishman & Volkman, 2018). This led to a bias in the Protein Data Bank towards globular, thermodynamically stable chains. However, the dynamic nature of protein NMR and the advancements of Cryo-EM as well as the redundancy among PDB entries (i.e. multiple entries that contain almost identical protein sequences) have offered valuable insights into protein conformational landscape (Hrabe et al., 2015; Monzon et al., 2016). A series of MD datasets simulating diverse protein structures have been released (Vander Meersche et al., 2023; Mirarchi et al., 2024), albeit they have limited sequence variability and only a fraction of the simulations provide valuable information into dynamics of large molecular conformational changes.
The large number of high-resolution deposited structures in the PDB allowed for the training of large geometric deep learning models that boosted the prediction accuracy of protein structure (Jumper et al., 2021; Kryshtafovych et al., 2019). Remarkable breakthroughs in protein backbone inverse folding (Dauparas et al., 2022) and de novo backbone generation have followed suit (Watson et al., 2023). ProteinMPNN (Dauparas et al., 2022) in particular has streamlined single-state protein inverse folding due to much smaller computational and time costs coupled with increased sequence recovery, when compared to physics-based methods (Liu & Kuhlman, 2006) that treat sequence design as an energy optimization problem. While these models have been successfully used for de novo protein design for improved function (Sumida et al., 2024), they mostly rely on enhancing protein stability and solubility rather than dealing with dynamics and interaction mechanisms.

### v2

@@
 We introduce DynamicMPNN, an inverse folding model explicitly trained to generate sequences compatible with multiple conformations through joint learning across conformational ensembles.
-Trained on 46,033 conformational pairs covering 75% of CATH superfamilies and evaluated using AlphaFold initial guess, DynamicMPNN outperforms ProteinMPNN by up to 13% on structure-normalized RMSD across our challenging multi-state protein benchmark.
+Trained on 46,033 conformational pairs covering 75% of CATH superfamilies and evaluated using Alphafold 3, DynamicMPNN outperforms ProteinMPNN by up to 25% on decoy-normalized RMSD and by 12% on sequence recovery across our challenging multi-state protein benchmark.
 1 Introduction
-The conformational diversity of proteins underlies biological functions such as enzyme catalysis, protein recognition, allostery, protein evolution modulation or human disease (Monzon et al., 2016).
-Of particular importance is the design of bioswitches, with huge implications in biotechnology, e.g.
-in creating artificial motors, signalling pathways, biosensors or drug delivery systems (Stein & Alexandrov, 2015; Praetorius et al., 2023).
-While most known switches undergo rearrangements in the context of a single fold (Ambroggio & Kuhlman, 2006a), the class of metamorphic proteins undergo changes in their secondary structure and fold (Fig.
+A commonly derived assumption from Anfinsen’s experiment is that proteins adopt only one native 3D structure, leading to the “one sequence, one structure, one function” canon.
+This view has been indirectly reinforced by the predominant use of X-Ray crystallography in experimental protein structure determination, which requires that proteins form a homogenous, diffractible crystal to be characterised (Dishman & Volkman, 2018).
+The large collection of static protein structures in the PDB has enabled the development of high-accuracy machine learning models for tasks such as structure prediction (Jumper et al., 2021; Kryshtafovych et al., 2019) and inverse folding (Dauparas et al., 2022; Hsu et al., 2022).
+Amongst contemporary inverse folding models, ProteinMPNN has been particularly widely adopted in applied protein design projects due to its low inference costs and robust experimental success rates (Dauparas et al., 2022; Watson et al., 2023; Goverde et al., 2024), outperforming traditional physics-based design methods on both fronts (Liu & Kuhlman, 2006).
+Although direct experimental characterisation of protein dynamics remains a challenge, the conformational diversity of proteins underlies crucial biological functions such as enzyme catalysis, protein-protein interactions, allostery, and human disease (Monzon et al., 2016).
+In applied protein design, bio-switches - proteins that switch between two structural states - are of particular importance, with key applications in engineering artificial bio-motors, signalling pathways, biosensors, or drug delivery systems (Stein & Alexandrov, 2015; Praetorius et al., 2023).
+While most known switches undergo rearrangements in the context of a single fold (Ambroggio & Kuhlman, 2006a), the class of metamorphic proteins undergo changes in both their secondary structure and fold (Fig.
 1a) and have been predicted to represent up to 4% of the PDB chains (Porter & Looger, 2018).
-These proteins are known to take two main functional states (Dishman & Volkman, 2018) and a finite number of conformations (see Discussion), thus this study is not aimed at the design of protein sequences adopting continuous conformational landscapes (e.g.
+These proteins typically adopt two main functional states (Dishman & Volkman, 2018) and a finite number of conformations (see Discussion).
+Beyond the world of switches, other dynamic proteins are characterised by continuous conformational landscapes (e.g.
 intrinsically disordered proteins (Tompa & Fuxreiter, 2008)).
-Multi-state protein design was first achieved through rational design and physics-based RosettaDesign (Liu & Kuhlman, 2006) of: metamorphic metal-binding peptides (Ambroggio & Kuhlman, 2006b; Cerasoli et al., 2005); closely related sequences that adopt diverging folds (Wei et al., 2020); and hinge proteins with binder-regulated equilibrium between states (Zhang et al., 2022; Quijano-Rubio et al., 2021; Praetorius et al., 2023).
-Remarkably, Praetorius et al.
-(2023) additionally employed ProteinMPNN Multi-state Design (ProteinMPNN-MSD) (Dauparas et al., 2022)—an inference strategy for extending ProteinMPNN to multiple states by averaging the logits of two independent single-state ProteinMPNN embeddings during the decoding step—to design proof-of-concept multi-helix hinge proteins.
-Praetorius et al.
-(2023) however report an overall very low experimental success rate: only 46 2-state hinge sequences, corresponding to roughly 0.002% of all designs, were successfully expressed in solution out of which only nine showed binding with the corresponding target peptide.
-This and the absence of any reported de novo designs for full natural multi-state backbones is proof that multi-state protein design lags behind single-state design (Appendix E) due to limited conformational datasets, weak benchmarks, and the poor record of folding models in predicting alternative states (Chakravarty et al., 2024) - which made for poor self-consistency filters.
+Multi-state protein design was first achieved through rational design and physics-based methods such as RosettaDesign (Liu & Kuhlman, 2006).
+Previous campaigns leveraging these methods have attempted to design metamorphic metal-binding peptides (Ambroggio & Kuhlman, 2006b; Cerasoli et al., 2005), closely related sequences that adopt diverging folds (Wei et al., 2020), and hinge proteins with binder-regulated thermodynamic equilibria, allowing the relative populations of different structural states to be modulated by exogenous proteins (Zhang et al., 2022; Quijano-Rubio et al., 2021; Praetorius et al., 2023).
+More recently, Praetorius et al.
+(2023) developed ProteinMPNN Multi-state Design (ProteinMPNN-MSD) (Dauparas et al., 2022): an inference strategy for extending ProteinMPNN inverse folding to multiple structural states by averaging the logits of two independent single-state ProteinMPNN embeddings during the decoding step.
+In a similar vein, ProteinGenerator (Lisanza et al., 2024) applies a sequence diffusion process conditioned on structural predictions by RosettaFold (Baek et al., 2021), and attempts to design sequences that adopt multiple states by averaging the sequence logits predicted by the model across distinct target structures.
+Still, significant limitations in current multi-state design pipelines remain.
+Designs generated from ProteinMPNN-MSD faced a rigorous selection process.
+From an initial pool of over two million computational designs, subsequent filtering and experimental validation yielded 46 soluble 2-state hinge sequences.
+Among these, nine successfully demonstrated binding to the target peptide (Praetorius et al., 2023).
+Likewise, the authors of ProteinGenerator reported a significantly lower in silico design success rate of 0.05% for their multi-state design task compared to rates of 2-10% observed for various single-state sequence design objectives using their framework (Lisanza et al., 2024).
+These observations - combined with the relative scarcity of published data on ML-driven multi-state de novo design campaigns - suggest that current ML methods for multi-state protein design have been significantly less successful than their single-state counterparts.
+We propose that this gap can be attributed to limited multi-conformational datasets, weak benchmarks, and the poor performance of folding models in predicting alternative states (Chakravarty et al., 2024) (which adversely affects their efficacy as self-consistency filters in protein design workflows).
 Our contributions. This paper introduces DynamicMPNN (Fig. 1b), a novel geometric deep learning-based pipeline for multi-state protein sequence design.
 DynamicMPNN is the first explicit multi-state inverse folding model for protein design.
-To train DynamicMPNN, we create a new ML-ready dataset of proteins with multiple conformations using the PDB and CoDNaS (Monzon et al., 2016) databases, and evaluate the method on 94 biologically relevant metamorphic, hinge, and transporter proteins.
-DynamicMPNN improves performance over ProteinMPNN (Dauparas et al., 2022) on AFIG by up to 13% on RMSD and 3% on pLDDT self-consistency values.
+To train DynamicMPNN, we create a new ML-ready dataset of proteins with multiple conformations using the PDB and CoDNaS (Monzon et al., 2016) databases, and evaluate the method on 96 biologically relevant metamorphic, hinge, and transporter proteins.
+We introduce a novel data processing pipeline and architecture able to handle sequence-aligned structure pairs with heterogeneous sequences, enabling us to leverage the conformational diversity across non-identical protein states.
+We propose a multi-state self-consistency metric and benchmark based on Alphafold3 (AF3) (Abramson et al., 2024) using target structures as templates.
+DynamicMPNN improves performance over ProteinMPNN (Dauparas et al., 2022) on AF3 by up to 25% on RMSD and 8% on pLDDT decoy normalized self-consistency values.
 2 The DynamicMPNN pipeline
 2.1 Protein multi-conformational dataset
-While over 750,000 individual protein chains (sequence-structure pairs) are available in the PDB, multi-conformational data is far more scarce with only 11,833 NMR-derived protein ensembles covering just 21% of CATH superfamilies.
-To overcome this limitation, we exploit the redundancy of chains with very high sequence similarity (95%) in the PDB to build a multi-conformational dataset of 46,033 conformer pairs that expands coverage to 75% of CATH superfamilies (Fig.
+While over 900,000 individual protein chains (sequence-structure pairs) are available in the PDB, multi-conformational data is at a first glance far more scarce with only roughly 12,000 NMR-derived protein ensembles covering just 21% of CATH superfamilies.
+To overcome this limitation, we exploit the sequence redundancy in the PDB: following Monzon et al.
+(2016) and Hrabe et al.
+(2015), we cluster all protein chain sequences with a very high sequence similarity threshold (95%) to build a multi-conformational dataset of 46,033 conformational clusters with at least 2 members - expanding coverage to 75% of CATH superfamilies (Fig.
 2a).
-For dataset splitting, we first curate a benchmark set from five previous studies of proteins with large conformational changes: 92 metamorphic proteins (Porter & Looger, 2018), 91 apo-holo proteins (Saldaño et al., 2022), the OC23/OC85 open-closed datasets, and 20 transporter proteins (Kalakoti & Wallner, 2025).
-The 94 highest RMSD pairs were assigned to the test set and the next 100 to the validation set.
-Training clusters were filtered to exclude any with TM-score (Zhang & Skolnick, 2004) to test/validation structures, preventing structural similarity leakage and yielding a final training set of 44,243 conformer pairs.
-See Appendix B for further details on dataset composition.
+The 95% similarity threshold accounts for soluble tags or point mutations/alterations performed in different experiments which are unlikely to affect the overall fold.
+Clusters contain highly unbalanced conformational populations (Fig.
+2c): a relatively small number of clusters contain thousands of sequences, while the majority contain only a few.
+This contributes to the relatively small intra-cluster RMSD in the majority of the clusters (Fig.
+2b), since many alternative protein states live in low populations.
+To maximise the conformational signal and minimize alignment artifacts, we built our dataset using only pairs of chains from one or two PDB entries that have the largest RMSD from each cluster (Fig.
+2c).
+Critically, this selection emphasizes our focus on proteins exploring a finite number of functional states, with most switching between just two conformations (Dishman & Volkman, 2018; Leaver-Fay et al., 2011; Alberstein et al., 2022).
+While we hope to explore design of more continuous conformational spaces in future work, we focus presently on training DynamicMPNN to design two-state proteins.
+Furthermore, molecular dynamics datasets were opted against due to the poor training signal given by limited sequence variability and scarcity of trajectories showing large conformational changes (Vander Meersche et al., 2023; Mirarchi et al., 2024), which normally happen across long biological timescales.
+At the other end of the spectrum, single-chain disordered protein simulations (Tesei et al., 2024) expose continuous conformational landscapes which are hard to discretize and have a high degree of stochasticity with a low signal-to-noise ratio rather than the ordered topologies of globular multi-state proteins (Tompa & Fuxreiter, 2008).
+For dataset splitting, we first curate a benchmark composed of five previous studies of proteins with large 2-state conformational changes: (1) 92 metamorphic proteins (Porter & Looger, 2018), (2) 91 apo-holo proteins (Saldaño et al., 2022), (3) the OC23 and OC85 open-closed datasets (Kalakoti & Wallner, 2025), and (4) 20 transporter proteins (Kalakoti & Wallner, 2025).
+The proteins with the highest inter-state RMSD were assigned to the test set (96 samples), while the rest were assigned to the validation set (100 samples).
+Training clusters were filtered to exclude any with TM-score (Zhang & Skolnick, 2004) and sequence similarity to test/validation structures, preventing structural similarity leakage and yielding a final training set of 44,243 conformer pairs.
+We additionally curate a set of single-state sequence-structure pairs from the 30% sequence similarity clusters not represented in our multi-conformational training dataset ().
+This augmentation strategy maximizes coverage of protein fold space while preserving the multi-conformational learning signal.
+See Appendix C for further details on dataset composition.
 2.2 DynamicMPNN for multi-state inverse folding
@@
 Extensions of these methods to multi-state design have thus far been limited to post-hoc aggregation of independent single-state predictions.
-Instead, DynamicMPNN learns the joint conditional distribution of directly through autoregressive sequence generation (see Appendix C), where represent distinct conformations; thus capturing sequence compatibility across all states simultaneously within a single model.
+We believe such methods favour logits which are highly biased towards one conformation, whose average over the 2 states is higher than the one of logits which are moderately valued across both states (Joshi et al., 2025).
+Instead, DynamicMPNN learns the joint conditional distribution of directly through autoregressive sequence generation, where represent distinct protein conformations encoded into a shared latent space; thus learning a sequence distribution that simultaneously satisfies multiple structural constraints.
+We decompose this joint conditional probability using the autoregressive factorization:
+where each factor represents the probability of selecting residue given the sequence prefix and the complete structural ensemble.
 Overall architecture.
-DynamicMPNN independently encodes each of the functional states of the protein, together with their interaction partners, into a shared latent feature space (Fig.
+DynamicMPNN independently encodes each of the functional states of a protein, together with their binding partners, into a shared latent feature space (Fig.
 1b).
-Embeddings of the target chains are then pooled across conformations to obtain a single embedding from which a sequence is auto-regressively generated.
-Our architecture is based on gRNAde (Joshi et al., 2025), a multi-state GNN for RNA inverse folding.
-For both the encoder and the decoder, we employ SE(3)-equivariant Geometric Vector Perception (Jing et al., 2021) layers which maintain computational efficiency through edge sparsity.
-Both encoder and decoder were assigned 8 GVP layers, following Hsu et al.
-(2022).
+Embeddings of the chains-to-be-designed chains are then pooled across conformations to obtain a single embedding from which a sequence is auto-regressively generated.
+Our architecture is built upon gRNAde (Joshi et al., 2025), a multi-state GNN for RNA inverse folding.
+For both encoder and decoder, we employ SE(3)-equivariant Geometric Vector Perceptron (Jing et al., 2021) layers which maintain computational efficiency through edge sparsity (k-NN edges with k=32).
+Within the GVP, the scalar features and vector features for each node (Duval et al., 2024):
+where are Geometric Vector Perceptrons, a generalization of MLPs to take tuples of scalar and vector features as input and apply -equivariant non-linear updates.
+The overall GNN encoder is -equivariant due to the use of reflection-sensitive input features (dihedral angles) combined with -equivariant GVP-GNN layers.
+Both encoder and decoder are assigned 8 GVP layers, following findings in Hsu et al.
+(2022) (2022).
 See Appendix D for further details.
+We present two different encoder architectures:
+DynamicMPNN: Independent encoder channels for each conformation, followed by Deep Set pooling (Zaheer et al., 2017) - it is invariant to conformation order and does not add extra parameters to our model.
+We note that while some more expressive pooling strategies have been shown to provide marginal performance improvements, they usually come at a great cost in efficiency (Joshi et al., 2025).
+Only node features are updated.
+DynamicMPNN + DSS: We implement cross-attention between the encoder channels after each layer using a Deep Symmetric Set (DSS) (Maron et al., 2020) module, which allows for richer inter-conformation interactions at the cost of increased computational complexity.
+In the scatter/gather DSS strategy, node embeddings of all design chains are averaged, passed through GVP layers, and added back to the residual features of each channel.
+Both node and edge features are updated.
+A major distinction from gRNAde is that DynamicMPNN models multi-chain complexes and can thus condition conformational changes on protein binders and oligomeric states.
+It may be possible to use this principle to engineer controllable protein conformational switches by tuning the free energy difference between folds and their binding interactions (Alberstein et al., 2022).
+However, this also introduces challenges in dealing with nonidentical sequences and missing residues across conformations, which we address as follows.
 Alignment and pooling.
-To deal with nonidentical sequences and missing residues during training, structure pairs are featurized and encoded independently, and subsequently aggregated based on pairwise sequence alignments (i.e., only nongap residues are taken into account when pooling).
+To deal with nonidentical sequences and missing residues during training, structure pairs are featurized and encoded independently, and subsequently aggregated based on pairwise sequence alignments (i.e.
+only nongap residues are taken into account when pooling).
 Multi-chain encoding and masking.
 Given that conformational shifts often depend on interactions, DynamicMPNN encodes the chemical environment for each conformational state (currently limited to proteins).
-During training, we expose the sequence information of binding partners in the encoding, only masking chains with 70% sequence similarity to the ground truth sequence of the chain of interest.
-Conformation order-invariant pooling.
-We employ Deep Set pooling (Zaheer et al., 2017) over the conformations as it is invariant to conformation order and does not add extra parameters to our model.
-We note that while some more expressive pooling strategies have been shown to provide marginal performance improvements, they usually come at a great cost in efficiency (Joshi et al., 2025).
-| RMSD | pLDDT | |||||||
-| Model | Best Paired | Best Paired | Best Paired | Best Single | All Avg. | Best Paired | Best Single | All Avg. |
-| (Å) | (Struct Norm) | (Decoy Norm) | (Å) | (Å) | ||||
-| DynamicMPNN | 13.43 | 3.86 | 0.58 | 12.76 | 17.39 | 58.65 | 59.31 | 50.22 |
-| ProteinMPNN-MSD | 14.76 | 4.47 | 0.65 | 13.99 | 17.24 | 55.89 | 57.74 | 50.18 |
-| Natural sequence | 16.85 | 5.51 | 0.76 | 16.85 | 16.85 | 48.51 | 48.51 | 48.51 |
+During training, we expose the sequence information of binding partners in the encoding, only masking chains with 70% sequence similarity to the ground truth sequence of the chains of interest.
 2.3 Multi-state design evaluation
-Following previous work (Wang et al., 2023), we evaluate the refoldability of generated sequences rather than sequence recovery (Appendix A).
-Existing refoldability methods compare target structures to single conformations predicted by folding models (e.g., AlphaFold2 (Jumper et al., 2021)).
-We argue that this approach is unsuitable for multi-state design since folding models typically predict one dominant state or interpolate between conformations rather than sampling the full conformational ensemble (Lane, 2023; Chakravarty et al., 2024; Saldaño et al., 2022).
-We propose using the AlphaFold initial guess (AFIG) framework (Roney & Ovchinnikov, 2022; Bennett et al., 2023), originally suggested to rank candidate protein structures, to evaluate multi-state designs (Praetorius et al., 2023).
-AFIG initializes the AF2 reference frames on the target backbone coordinates during inference to bias the model towards the target conformation.
-The similarity (C-RMSD) between predicted and target structures, along with AF2 confidence scores, serves as a proxy for the likelihood that the designed sequence will fold into the target structure.
+Following previous work Wang et al.
+(2023), we evaluate the refoldability of generated sequences rather than sequence recovery (Appendix B).
+Existing refoldability methods compare target structures to single conformations predicted by folding models (e.g., AlphaFold2 Jumper et al.
+(2021)).
+We argue that this approach is unsuitable for multi-state design since folding models typically predict one dominant state or interpolate between conformations rather than sampling the full conformational ensemble Lane (2023); Chakravarty et al.
+(2024); Saldaño et al.
+(2022).
+Furthermore, as mentioned in Section 2.1, recent DL-based protein folding models have been shown not to learn protein folding physics, but rather to compress the protein space information while preserving lower complexity motifs such as sequence segment pairing (Outeiral et al., 2022; Zhang et al., 2024; Chakravarty et al., 2024).
+We propose using a template-based Alphafold 3 (AF3) framework Abramson et al.
+(2024): for each de novo sequence, we perform 2 AF3 runs, one with each conformational state given as template to moderately bias the model towards the target conformation.
+The similarity - C-RMSD or TM-score (Zhang & Skolnick, 2004) - between predicted structures and the templates, along with AF3 confidence scores, serve as a proxy for the likelihood that the designed sequence will fold into the target structure.
 This framework naturally enables evaluation of refoldability across multiple conformations.
-Formally, for a protein with conformational states and designed sequence , we define the AFIG RMSD for each target conformation as:
-where is the structure predicted by AlphaFold2 for sequence when backbone frames are initialized on coordinates of structure .
-We define two normalization criteria to contextualize observed deviations:
-Structure normalization (Struct Norm): We normalize AFIG RMSD by the maximum RMSD between target conformations to capture task difficulty; i.e.
-designing proteins with large conformational shifts is inherently more difficult than small rearrangements:
-Decoy normalization (Decoy Norm): We initialize AF2 with structurally dissimilar decoy structures () using the same sequences designed and measure the resulting deviations.
+Formally, for a protein with conformational states and designed sequence , we define the AF3 RMSD for each target conformation as:
+where is the structure predicted by AlphaFold3 for sequence when a template of is provided.
+To account for the structural bias induced by the template, we define a normalization strategy to contextualize observed deviations:
+Decoy normalization (Decoy Norm): We provide AF3 structurally dissimilar decoy structures as templates () using the same sequences designed and measure the resulting deviations.
 This control assesses whether sequences fold specifically into their targets or may fold equally well into arbitrary structures:
 where is a decoy that is structurally dissimilar to .
-Additionally, we measure pLDDT confidence scores to evaluate AFIG fold uncertainty.
+Additionally, we measure pLDDT confidence scores to evaluate AF3 fold uncertainty.
 High RMSD with low pLDDT indicates poor template matching, while low RMSD with high pLDDT suggests a successful design.
+| Raw Metrics | Decoy-Normalized Metrics | |||||
+| Model Variant | pLDDT | RMSD (Å) | TM-score | pLDDT | RMSD | TM-score |
+| Combined Training | 79.28 (9.28) | 4.37 (4.76) | 0.774 (0.246) | 1.513 (0.503) | 0.200 (0.183) | 7.396 (2.192) |
+| Combined Training + DSS | 78.87 (8.82) | 4.38 (4.47) | 0.769 (0.239) | 1.532 (0.524) | 0.200 (0.164) | 7.384 (2.304) |
+| Single Training | 62.59 (16.84) | 13.48 (13.02) | 0.503 (0.350) | 1.466 (0.537) | 0.479 (0.339) | 5.014 (3.068) |
 3 Results and Discussion
+| Model Variant | Sequence Recovery (%) |
+| Combined Pretraining + Multi Finetuning | 42.7 (8.8) |
+| Single Pretraining + Multi Finetuning | 42.1 (8.3) |
+| Combined Pretraining + Multi Finetuning + DSS | 41.9 (8.6) |
+| Combined Training | 41.0 (8.5) |
+| Single Pretraining + Multi Finetuning + DSS | 40.3 (8.2) |
+| Combined Training + DSS | 38.8 (7.8) |
+| ProteinMPNN MSD* | 38.0 (11.0) |
+| Single Training | 27.1 (9.4) |
+| Single Training + DSS | 26.2 (8.5) |
+| Raw Metrics | Decoy-Normalized Metrics | |||||
+| Model | pLDDT | RMSD (Å) | TM-score | pLDDT | RMSD | TM-score |
+| Combined Training | 79.23 (7.79) | 4.30 (3.50) | 0.748 (0.233) | 1.401 (0.411) | 0.206 (0.170) | 7.093 (2.034) |
+| Combined Training + DSS | 78.90 (7.68) | 4.56 (3.52) | 0.737 (0.225) | 1.404 (0.425) | 0.217 (0.170) | 7.090 (2.384) |
+| Single Training | 63.10 (15.65) | 13.21 (12.18) | 0.477 (0.331) | 1.364 (0.458) | 0.508 (0.337) | 4.527 (2.777) |
+| ProteinMPNN MSD* | 74.29 (11.37) | 5.84 (4.56) | 0.689 (0.244) | 1.416 (0.389) | 0.275 (0.219) | 6.533 (2.325) |
 Setup.
-DynamicMPNN was trained for 50 epochs, and an AFIG evaluation on the validation set was performed every third epoch to select the best-performing model.
-Then, DynamicMPNN and ProteinMPNN (using Multi-state Design inference strategy) were used to sample 16 sequences for the benchmark set of 94 paired conformations, which were each run through the AFIG pipeline separately against both target states.
-Aggregated evaluation metrics were computed as follows.
-(Best Single) The minimum RMSD and maximum pLDDT values across all sets of 16 sequences are selected from among both states.
-(Best Paired) The RMSD and pLDDT values are averaged over the 2 states for each sequence, then the best RMSD and pLDDT across all 16 averaged values are selected.
-Struct Norm and Decoy Norm RMSD values are then calculated for each protein.
-(All Avg.) The RMSD and pLDDT values are averaged across all 16 sequences and both states.
-The natural sequences were processed identically and resulting metrics were averaged over the benchmark set and reported in Table 1.
-DynamicMPNN outperforms ProteinMPNN.
-We see this across all metrics, with the largest gains in the ”Best Paired” metrics of up to 13% - arguably the most important set of metrics for experimental design, as only the best sequences are selected for in vitro validation.
-Visual structural inspection of both states (Fig.
-5) confirms successful design of the metamorphic Switch Arc protein by DynamicMPNN, compared to failure to recapitulate the central beta sheet fold by the best scoring ProteinMPNN sequence.
-DynamicMPNN also outperforms ProteinMPNN in the ”Best Single” metrics, which is surprising since ProteinMPNN was trained on single-state design tasks.
-It should be noted that ProteinMPNN training dataset contains proteins within the sequence clusters of 84 out of 94 benchmark proteins.
-This represents a substantial data leakage advantage for ProteinMPNN, while DynamicMPNN was specifically trained to exclude any sequence similarity to the test set.
-We plan on training a single-state version of DynamicMPNN to more fairly demonstrate the advantage of the multi-state design approach.
-While DynamicMPNN AFIG RMSD values are still high, it is important to contextualize these results within our deliberately challenging benchmark: our test set comprises proteins with the largest documented conformational changes in the known proteome.
+We evaluate several training strategies and model architectures to assess the impact on multi-state learning.
+Single Training models follow the standard single state inverse folding approach, learning exclusively from single-state sequence structure pairs, similarly to ProteinMPNN Dauparas et al.
+(2022).
+In Combined Training, models are presented with both single and multi-state training samples.
+We also explore training over multiple stages, where single-state data is provided during pretraining to teach general principles of protein geometry from abundant single-state structures before specializing on multi-state conformational pairs (Multi Finetuning).
+For each model, the training checkpoint with the highest sequence recovery on the multi-state validation set was selected for evaluation.
+All DynamicMPNN versions were trained on either 8 A100-80GB or 8 H100-80GB GPUs using a batch size of 32 and Adam (Kingma, 2014) optimizer with learning rate .
+Training for each stage was run until the respective validation sequence recovery (on single-state validation dataset for single training runs and multi-state for multi-state runs) plateaued or began to decline, typically after 20-50 epochs.
+Then, DynamicMPNN and ProteinMPNN (using Multi-state Design inference strategy) were used to sample 16 sequences for the benchmark test set of 96 paired conformations, which were each run through the AF3 pipeline separately against both target states.
+Refoldability was evaluated for designed sequences of the model checkpoints with the highest sequence recovery scores by predicting the structure using AF3 with target and decoy structures as templates.
+Aggregated evaluation metrics were computed by averaging across all 16 sequences and both states.
+DynamicMPNN outperforms existing benchmarks across multiple evaluation metrics.
+Our best-performing model achieves substantial improvements over baseline methods: a 25% reduction in decoy-normalized RMSD (Tab.
+3) and a 12% improvement in sequence recovery (Fig.3; Tab.
+2) compared to ProteinMPNN Multi-State Design (MSD).
+This performance gain is particularly noteworthy considering that ProteinMPNN’s training dataset contains proteins within 30% sequence similarity clusters of 91 out of 96 benchmark proteins, making our improvement on this established baseline significant.
+We leave the retraining of ProteinMPNN on a rigorous train-test split for future work.
+However, a more rigorous comparison involves our single-state trained model, which adheres to the same training-test split and thus eliminates potential data leakage concerns.
+Against this baseline, DynamicMPNN demonstrates substantial improvements across all evaluation metrics, providing clear evidence that explicit multi-state training is superior to single-state approaches for multi-conformational protein design tasks (Tab.
+2).
+Our single-state trained model performs considerably worse than both DynamicMPNN variants and ProteinMPNN MSD, achieving only 63.10 pLDDT compared to 79.23 for our best model, and showing nearly three-fold higher RMSD values (13.21 Å vs 4.30 Å; Tab.
+1).
+Some likely reasons to explain the under performance of our model checkpoint trained exclusively on single-state data is the aforementioned unfair train splitting of ProteinMPNN.
+Combined and multi-state finetuning training strategies far outperform the single-state only training Models trained only on single-state data (Single Training) perform poorly on multi-state design as they are unable to create and decode a meaningful latent representation of the conformational changes.
+The Combined Training approach, which exposes models to both single-state and multi-state examples during training, achieves the optimal balance and consistently outperforms other training strategies.
+While DynamicMPNN AF3 RMSD values are still high, it is important to contextualize these results within our deliberately challenging benchmark: our test set comprises proteins with the largest documented conformational changes in the known proteome.
 This can be seen as an extrapolation beyond the overall modest conformational changes observed in most PDB structures and hence our training data.
-Natural sequence vs.
-multi-state designs.
-While the best-of-16 sequence from DynamicMPNN and ProteinMPNN (Dauparas et al., 2022) outperform natural sequences, the average metrics are slightly worse than the natural sequences (Fig.
-LABEL:fig:supplementary_plots).
-Dauparas et al.
-(2022) have presented a similar finding with respect to single-state protein design, where the ProteinMPNN designed sequences showed improved AlphaFold refoldability over the natural sequences.
-A possible explanation is that while both inverse folding models are trained mostly to optimize the sequence for structure stability, evolution has undergone a multi-objective optimization for different functions, which might have prevented it from reaching the energy minima for those structures.
-Additionally, due to graph sparsity, DynamicMPNN and ProteinMPNN emphasize local interactions, making designed sequences more foldable by AFIG - since AFIG does not use any MSA, it poorly captures long-range interactions present in natural sequences (Zhang et al., 2024).
 4 Conclusion
-We present DynamicMPNN, the first explicit multi-state inverse folding model, achieving up to 13% improvement over ProteinMPNN on our multi-state benchmark.
+We present DynamicMPNN, the first explicit multi-state inverse folding model, achieving up to 25% improvement over ProteinMPNN on our multi-state benchmark.
 By jointly learning across conformational ensembles rather than aggregating single-state predictions, DynamicMPNN captures sequence constraints required for multiple functional conformations.
 This opens possibilities for engineering synthetic bioswitches, allosteric regulators, and molecular machines.
-Software and Data
-We provide the code for our model in the following repository: github.com/Alex-Abrudan/DynamicMPNN.
-Acknowledgements
-The authors would like to thank Adam Broerman, Rohith Krishna, Kai Yi, Simon Mathis, Hannah Han, Vladimir Radenkovic, Gustavo Parisi, Sebastian Ahnert, and Pietro Liò for insightful comments and discussions.
-CKJ was supported by the A*STAR Singapore National Science Scholarship (PhD).
-AK was supported by the Graduate School of Quantitative Biosciences Munich (QBM).
-Impact Statement
-This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
-Appendix A Supplementary Results
-A.1 Additional Performance Analysis
-For our statistical analysis, we first evaluated the normality of the paired differences between model performances (RMSD values) using the Shapiro-Wilk test.
-The test results (W = 0.913, p 0.001) strongly rejected the null hypothesis of normality for the distribution of differences.
-Therefore, we selected the Wilcoxon signed-rank test as the appropriate non-parametric alternative to the paired t-test for assessing statistical significance between model performances.
-This test makes no assumptions about the underlying distribution shape, making it robust for our analysis of paired protein structure comparisons across the different design approaches.
-Fig.
-LABEL:fig:supplementary_plots (a), (e) shows how DynamicMPNN maintains its superiority for the single-state best sequence design for the RMSD and pLDDT metrics.
-However, averaging the performance of all designed sequences erases the advantage of DynamicMPNN over ProteinMPNN and natural sequences (Fig.
-LABEL:fig:supplementary_plots (b), (c), (d), (f)).
-While ProteinMPNN outperforms DynamicMPNN on sequence recovery and perplexity (Table 2), Wang et al.
-(2023) argue that refoldability - emphasised in the present work - is more direct, structure-grounded, and superior to sequence recovery and perplexity when it comes to evaluating inverse folding models.
-| Model | Sequence Recovery | Perplexity |
-| DynamicMPNN | 0.29 | 10.2 |
-| ProteinMPNN MSD | 0.38 | 5.12 |
-Appendix B Dataset Details
+It is unclear if the presented one-size-fits-all approach to multi-state design will be effective experimentally, or if specialized models for different classes of conformational changes depending on their thermodynamic complexity will be beneficial.
+Appendix A Appendix
+A.1 LLM acknowledgement
+The authors acknowledge that they have used LLMs in the process of paper writing for style suggestions and proofreading.
+Appendix B Supplementary Results
+B.1 DynamicMPNN preliminary designs
+Appendix C Dataset Details
 To construct the dataset, we obtained 46,033 Multiple Sequence Alignment (MSA) clusters at local sequence similarity from the latest version of CoDNaS (v2025) (Monzon et al., 2016), including NMR model structures.
 Importantly, the CoDNaS dataset prevents potential errors created by mixing different homologues in the same cluster by enforcing the same UniProt ID for all custer members.
-The 95% similarity threshold accounts for soluble tags or point mutations/alterations performed in different experiments so that all available conformations of a protein are included - as different experimental conditions and sequence variations can reveal distinct thermodynamic states of the same protein (Best et al., 2006).
+All available conformations of a protein are included - as different experimental conditions and sequence variations can reveal distinct thermodynamic states of the same protein (Best et al., 2006).
 While clusters contain varying numbers of conformations (Fig.
@@
 2c).
-Since the proteins targeted by this project explore two functional states, this choice aims to minimize alignment artifacts and maximize conformational signal during training - chosen pairs represent the most distinct conformational states.
+Chosen pairs represent the most distinct conformational states.
 We will include conformational information of the whole cluster in future work.
-Appendix C Multi-conformational inverse folding
-Given a set of target conformations we seek to model the conditional probability distribution over amino acid sequences that can adopt all specified structures.
-The challenge lies in learning a sequence distribution that simultaneously satisfies multiple structural constraints.
-We decompose this joint conditional probability using the autoregressive factorization:
-where each factor represents the probability of selecting residue given the sequence prefix and the complete structural ensemble.
+While other inverse folding models (Hsu et al., 2022) saw improved performance when trained on orders of magnitude of more protein structures from AFDB, previous studies have found that the majority of high-confidence structures in AFDB map to known CATH superfamilies (Bordin et al., 2023), and that AlphaFold struggles in predicting alternative states (Chakravarty et al., 2024).
+We therefore decided against including AFDB structures in our training set.
 Appendix D Model Details
@@
 After using message passing layers that are conformation order-equivariant, we add a conformation order-invariant head, which performs average pooling across the conformation channel of the scalar and vector feature tensors, similar to Joshi et al.
-(2025): and to and , where is the sequence length, is the number of backbones, is the number of scalar features, and is the number of vector features.
+(2025) (2025): and to and , where is the sequence length, is the number of backbones, is the number of scalar features, and is the number of vector features.
 The only pooling strategy used in this work is the pooling of the maximum RMSD pair of chains - therefore - although more pooling strategies for homo-oligomers can be used, such as equal averaging of all chains to be inverse folded in the selected PDB entries.
-Appendix E Further Discussion
-E.1 Motivation
-A commonly derived assumption from Anfinsen’s experiment is that proteins adopt only one native 3D structure, leading to the “one sequence, one structure, one function” canon.
-This view has been indirectly reinforced by the predominant use of X-Ray crystallography in experimental protein structure determination - it requires the protein to form a single, diffractible crystal (Dishman & Volkman, 2018).
-This led to a bias in the Protein Data Bank towards globular, thermodynamically stable chains.
-However, the dynamic nature of protein NMR and the advancements of Cryo-EM as well as the redundancy among PDB entries (i.e.
-multiple entries that contain almost identical protein sequences) have offered valuable insights into protein conformational landscape (Hrabe et al., 2015; Monzon et al., 2016).
-A series of MD datasets simulating diverse protein structures have been released (Vander Meersche et al., 2023; Mirarchi et al., 2024), albeit they have limited sequence variability and only a fraction of the simulations provide valuable information into dynamics of large molecular conformational changes.
-The large number of high-resolution deposited structures in the PDB allowed for the training of large geometric deep learning models that boosted the prediction accuracy of protein structure (Jumper et al., 2021; Kryshtafovych et al., 2019).
-Remarkable breakthroughs in protein backbone inverse folding (Dauparas et al., 2022) and de novo backbone generation have followed suit (Watson et al., 2023).
-ProteinMPNN (Dauparas et al., 2022) in particular has streamlined single-state protein inverse folding due to much smaller computational and time costs coupled with increased sequence recovery, when compared to physics-based methods (Liu & Kuhlman, 2006) that treat sequence design as an energy optimization problem.
-While these models have been successfully used for de novo protein design for improved function (Sumida et al., 2024), they mostly rely on enhancing protein stability and solubility rather than dealing with dynamics and interaction mechanisms.